OpenStudy (anonymous):

(1/8x^4+2/7x^3+7/8x^2+7)+(-5/8x^4+1/8x^2-7)

OpenStudy (anonymous):

\[-1/(2x^4)+2/(7x^3)+1/x^2\]

OpenStudy (anonymous):

ok im lost on this

OpenStudy (turingtest):

Just to confirm the equation is: \[(1/8x^4+2/7x^3+7/8x^2+7)+(-5/8x^4+1/8x^2-7)\] ???

OpenStudy (anonymous):

yea but i posted a different question after that. you are more than welcome to do that one if you like. its the 1/16x^4 one

OpenStudy (anonymous):

I'm guessing it's this \[(\frac{1}{8}x^{4} + \frac{2}{7}x^{3}+\frac{7}{8}x^{2}+7)+(-\frac{5}{8}x^{4}+\frac{1}{8}x^{2}-7)\]